The equation of a transverse sinusoidal wave is given by: y = 2x10^-3 sin (18x - 600t +30⁰) m. Find the:


(a) amplitude of the wave


(b) the wavelength


(c) the frequency


(d) the wave speed, and


(e) the displacement at position 0m and time 0s


(f) the maximum transverse particle speed.​

1
Expert's answer
2021-12-12T16:41:52-0500

(a) "A=0.002\ (m)"


(b) "\lambda=2\pi/k=2\cdot3.14/18=0.35\ (m)"


(c) "\nu=\omega/(2\pi)=600/(2\cdot3.14)=95.5\ (Hz)"


(d) "v=\omega/k=600/18=33.3\ (m/s)"


(e) "y=0.002\cdot\sin(0-0+30)=0.001\ (m)"


(f) "v_{max}=A\omega=0.002\cdot600=1.2\ (m/s)"